OpenStudy (anonymous):

find the least number which on decreasing by 20 is exactly divisible by 18,21,28,30

OpenStudy (anonymous):

okay, so find the LCM of 18 21 28 and 30. 18 = 3*3*2 21 = 3*7 28= 2* 2 *7 30 = 2* 3 * 5 therefore LCM is 3*3*2*7*2*5 = 1260. When you decrease 1280 by 20, you get 1260, the LCM.

OpenStudy (anonymous):

Do you get it?